You are considering investing in a mutual fund’s A share class that has a 3% sales charge (front-end load) and annual expense ratio of 1.0%. Alternatively, you might invest in that fund’s C shares that have no load and an annual expense ratio of 1.5%. Assume the fund’s assets return 9% annually.
a. [3 points] For a 2-year holding period, which share class will you prefer?

b. [3 points] For a 50-year holding period, which share class will you prefer?

c. [5 points] After how many years will the two mutual fund share classes result in the same future wealth amounts?

d. [2 points] How does your answer to part c. change is the fund’s assets return 14% per year? e. [2 points] What accounts for the greater expenses of the C-shares, compared to the Ashares?
Answer #1

a) Value of the share class can be calculated as below

FV = PV x (1 + r)^n

Assume you invest $100 in each of them,

Value in A, FV = 100 x (1 - 3%) x (1 + 9% - 1%)^2 = $113.14

Value in C, FV = 100 x (1 + 9% - 1.5%)^2 = $115.56

b) In A, FV = 100 x (1 - 3%) x (1 + 9% - 1%)^50 = $4,549.46

in C, FV = 100 x (1 + 9% - 1.5%)^50 = $3,718.97

c) In 7 years, value of both funds = $166

d) In 7 years, value of both funds = $228